OpenStudy (anonymous):

Carolyn is building a pen for her dog. The perimeter of the pen is 52 ft a) Explain how you can use a proportion to convert the perimeter into yards and feet. b) The fencing material is sold by the yard. It costs 10.99 per yard. What is the cost of material before taxes Please show your work

OpenStudy (anonymous):

NICE PIC

OpenStudy (anonymous):

What shape is it.

OpenStudy (anonymous):

square

OpenStudy (anonymous):

a) wouldnt you just divide it by 4? b)52/3=17 1/3 17 1/3*10.99= 190.493333333...(you get the point)

OpenStudy (anonymous):

kkk thanx